The force of friction between an object and the surface upon which it is sliding is 15N. The mass of the object is 20kg. What is the coefficient of kinetic friction?





1
Expert's answer
2020-12-23T07:34:54-0500

"F_k=15N"

"F= m*g= 20*9.8=196N"

"\text{The frictional force between two surfaces after sliding begins is the}"

"\text{product of the kinetic friction coefficient and the normal force:}"

"F_k=\mu_k*F"

"\mu_k=\frac{F_k}{F}=\frac{15}{196}\approx0.077"

Answer:"\mu_k\approx0.077"